# Stereo Anything: Unifying Stereo Matching with Large-Scale Mixed Data

### v1

@@ Abstract
Stereo matching has been a pivotal component in 3D vision, aiming to find corresponding points between pairs of stereo images to recover depth information. In this work, we introduce StereoAnything, a highly practical solution for robust stereo matching. Rather than focusing on a specialized model, our goal is to develop a versatile foundational model capable of handling stereo images across diverse environments. To this end, we scale up the dataset by collecting labeled stereo images and generating synthetic stereo pairs from unlabeled monocular images. To further enrich the model’s ability to generalize across different conditions, we introduce a novel synthetic dataset that complements existing data by adding variability in baselines, camera angles, and scene types. We extensively evaluate the zero-shot capabilities of our model on five public datasets, showcasing its impressive ability to generalize to new, unseen data (Figure 1). Code will be available at https://github.com/XiandaGuo/OpenStereo.
@@
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding authors: long.chen@ia.ac.cn
@@ 2.3 Multi-Dataset Training
To tackle the challenge of generalization, some researchers have increasingly explored the use of multi-dataset training strategies. By combining multiple datasets, models can be trained to capture diverse scene characteristics better, making them more robust. Although well-established in fields like object recognition [43], segmentation [21], and monocular depth estimation [31, 49, 19, 50], the use of this strategy in stereo matching remains relatively underexplored.
@@ 3 Existing Datasets
To enhance the robustness of the stereo-matching model, we incorporate a large-scale collection of 20.27 million unlabeled monocular images sourced from four public datasets. A comprehensive overview of these datasets is provided in Table 2. These will be used to generate further, synthetic stereo pairs during training.
@@
DrivingStereo [48] is a large-scale stereo dataset which includes over 180K images covering a diverse set of real-world driving scenarios. Sparse ground-truth annotations are obtained from LiDAR scans, being post-processed through a deep neural network.
@@ 6.3 Comparison with other SOTA methods
As shown in Table 3, we compare our best-performing model with several SOTA stereo-matching methods. StereoAnything achieves the lowest error rates on the KITTI 2012 [11], KITTI 2015 [29], and ETH3D [35] datasets, outperforming previous methods. In Table 4, we further present an ablation study that highlights the significant impact of our proposed training strategy, which applied to different stereo backbones yields significant improvement in performance across all datasets. Our training approach consistently delivers significant performance improvements across all datasets, showcasing its versatility and effectiveness. The most striking improvement is seen in the NMRFStereo-SwinT [13] variant shows particularly significant improvement, with the mean error reduced from 18.11 to 5.01. This notable gain is attributed to the increased data scale, which allows the SwinTransformer [26] backbone to unlock its full potential.
@@ 6.4 Ablation Studies
As shown in Table 8, we first validate the impact of using different monocular depth estimation models to generate pseudo stereo data on the BDD100k dataset. Fine-tuning on the stereo datasets synthesized using MiDas [31], DepthAnytingV1 [49] (DAv1), and DepthAnytingV2 [50] (DAv2) leads to competitive results, demonstrating the effectiveness of using monocular depth models for pseudo stereo data generation. Specifically, the synthesized data from DAv2 [50] outperforms other monocular methods across all benchmarks.
Then, we evaluate the impact of using different monocular datasets on the performance of the pseudo stereo data generation, based on the DAv2 [50] model. As shown in Table 9, the Google Landmarks [45] achieves the best performance overall, with a mean metric of 5.91. The slightly worse performance of ImageNet21K [32] (21K) compared to the GL [45] dataset can be attributed to the lower resolution of ImageNet21K images, which is smaller than our model’s input size. During the process, we resized the ImageNet21K [32] images to fit the input size of our model before running the monocular model to estimate depth, followed by generating the pseudo stereo pairs. This resizing process might have led to a loss of fine details, thereby affecting the quality of the generated depth and resulting in slightly lower performance compared to GL [45].

### v2

@@ Abstract
Stereo matching has been a pivotal component in 3D vision, aiming to find corresponding points between pairs of stereo images to recover depth information. In this work, we introduce StereoAnything, a highly practical solution for robust stereo matching. Rather than focusing on a specialized model, our goal is to develop a versatile foundational model capable of handling stereo images across diverse environments. To this end, we scale up the dataset by collecting labeled stereo images and generating synthetic stereo pairs from unlabeled monocular images. To further enrich the model’s ability to generalize across different conditions, we introduce a novel synthetic dataset that complements existing data by adding variability in baselines, camera angles, and scene types. We extensively evaluate the zero-shot capabilities of our model on five public datasets, showcasing its impressive ability to generalize to new, unseen data (Figure 1).
@@
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding authors
@@ 2.3 Multi-Dataset Training
To tackle the challenge of generalization, some researchers have increasingly explored multi-dataset training strategies. By combining multiple datasets, models can be trained to capture diverse scene characteristics better, making them more robust. Although well-established in fields like object recognition [43], segmentation [21], and monocular depth estimation [31, 49, 19, 50], the use of this strategy in stereo matching remains relatively underexplored.
@@ 3 Existing Datasets
To enhance the robustness of the stereo-matching model, we incorporate a large-scale collection of 30.07 million unlabeled monocular images sourced from four public datasets. A comprehensive overview of these datasets is provided in Table 2. These will be used to generate further, synthetic stereo pairs during training.
@@
DrivingStereo [48] is a large-scale stereo dataset that includes over 180K images covering a diverse set of real-world driving scenarios. Sparse ground-truth annotations are obtained from LiDAR scans, being post-processed through a deep neural network.
@@ 6.3 Comparison with other SOTA methods
As shown in Table 3, we compare our best-performing model with several SOTA stereo-matching methods. StereoAnything achieves the lowest error rates on the KITTI 2012 [11], KITTI 2015 [29], and ETH3D [35] datasets, outperforming previous methods. In Table 4, we further present an ablation study that highlights the significant impact of our proposed training strategy, which applied to different stereo backbones yields significant improvement in performance across all datasets. Our training approach consistently delivers significant performance improvements across all datasets, showcasing its versatility and effectiveness. The most striking improvement is seen in the NMRFStereo-SwinT [13] variant, which shows particularly significant improvement, with the mean error reduced from 18.11 to 5.01. This notable gain is attributed to the increased data scale, which allows the SwinTransformer [26] backbone to unlock its full potential.
@@ 6.4 Ablation Studies
As shown in Table 8, we first validate the impact of using different monocular depth estimation models to generate pseudo stereo data on the BDD100k [52] dataset. Fine-tuning on the stereo datasets synthesized using MiDas [31], DepthAnytingV1 [49] (DAv1), and DepthAnytingV2 [50] (DAv2) leads to competitive results, demonstrating the effectiveness of using monocular depth models for pseudo stereo data generation. Specifically, the synthesized data from DAv2 [50] outperforms other monocular methods across all benchmarks.
Then, we evaluate the impact of using different monocular datasets on the performance of the pseudo stereo data generation, based on the DAv2 [50] model. As shown in Table 9, the Google Landmarks [45] achieves the best performance overall, with a mean metric of 5.91. The slightly worse performance of ImageNet21K [32] (21K) and LSUN [51] compared to the GL [45] dataset can be attributed to the lower resolution, which is smaller than our model’s input size. During the process, we resized the images of 21K [32] and LSUN [51] to fit the input size of our model before running the monocular model to estimate depth, followed by generating the pseudo stereo pairs. This resizing process might have led to a loss of fine details, thereby affecting the quality of the generated depth and resulting in slightly lower performance compared to GL [45].